This is so complicated the sum of two numbers is 77 The second is 5 times more than 3 times the first.What are the two numbers? I do not understand
Answer by checkley71(8403)

77=X+Y OR
THE FIRST NUMBER IS X & THE SECOND (Y) IS (3X+5) SO WE SUBSTITUTE THESE VALUES IN THE FIRST EQUATION AND WE GET
77=X+(3X+5)
77=X+3X+5
4X=77-5
4X=72
X=72/4
X=18 FOR THE FIRST NUMBER AND
3*18+5=54+5=59 FOR THE SECOND NUMBER
PROOF
18+59=77
77=77